OpenStudy (anonymous):

hhFor which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. x/9 - 8 = 9 D. 4x + 5 = 41

OpenStudy (anonymous):

a

Directrix (directrix):

If x = 9, then 4(9) + 5 = 36 + 5 = 41. D

OpenStudy (anonymous):

soryy Directrix is right, it would be a if they asked for -9 sorry

Directrix (directrix):

Option A does not work: 3x – 7 = –34 3(9) - 7 = 27 - 7 = 20 which is not -34.

OpenStudy (anonymous):

thanks

Directrix (directrix):

Happy to help.

OpenStudy (anonymous):

:)